OpenStudy (anonymous):

Please check if my answer's right??? √3 (√6 + √3) + (√8 - 5) Answer: √18 + √8-5

OpenStudy (anonymous):

sqrt(3) sqrt(6) = sqrt(18) = 3sqrt(2) sqrt(3) sqrt(3) = 3 sqrt8=2sqrt2 sqrt(3)(sqrt(6)+sqrt(3)) +(sqrt(8)-5) =3sqrt2+3+2sqrt2 -5 =5sqrt2 -2